THE ROTATIONAL SPECTRUM OF THE H_{2}-OCS VAN DER WAALS COMPLEX

\underline{ZHENHONG YU} and WILLIAM KLEMPERER, Department of Chemistry and Chemical Biology, Harvard University, Cambridge, MA 02138; KELLY HIGGINS, Department of Chemistry, University of Minnesota, Minneapolis, MN 55455.

The microwave spectrum of the H_{2}-OCS complex has been studied using molecular beam electric resonance optothermal spectrometer with 20% 40:1 H_{2}/OCS gas mixture in 5 atm He. This investigation is stimulated by the recent He droplet study of this species^{a}. The H_{2}-OCS complex has T-shaped global minimum according to a high-level ab initio potential surface calculation. We have performed bound state calculations with bound program^{b} to obtain the rotational transitions of the ground state for para-H_{2}-OCS complex. Several spectral transitions of ^{32}S and ^{34}S isomers have been measured at frequencies close to the prediction in the frequency region of 9-31 GHz. For example, two observed peaks at 10218.3 and 20385.7 MHz from H_{2}-OC^{32}S complex have been compared with the predicted 1_{01}←0_{00} and 2_{02}←1_{01} transitions at 10594.7 and 21110.2 MHz, respectively. However a number of transitions predicted were not observed although the analogous transitions in the He-OCS complex have been readily seen^{c}. In addition, small doublet splitting of near 50 KHz is observed. Since ortho-H_{2} contributes to about 75 percent abundance of normal H_{2} at room temperature and might bind stronger with OCS molecule than para-H_{2}, we believe that what we observed in this study may belongs to ortho-H_{2}-OCS complex. Our current experimental measurement with pure para-H_{2} would provide detailed results for rotational transitions of the ground state for H_{2}-OCS complex in the near future.

aS. Grebenev, B. G. Sartakov, J. P. Toennies and A. F. Vilesov, J. Chem. Phys. \underline{114}(2), 617, 2001.

bJ. M. Hutson, Bound Computer Program, Version5, 1993 .

cKelly Higgins and William Klemperer, J. Chem. Phys. \underline{110}(3), 1383, 1999.